OpenStudy (anonymous):

solve this equation by using systems of elimination please! -x+5y=13 x-y=15

OpenStudy (anonymous):

elimination is the process of stacking them together and adding them (or subtracting them) so that variable(s) are eliminated. try adding them together.

OpenStudy (anonymous):

Eliminating the x's you're left with \[4y = 28\]\[y = 7\] Plugging 7 in one of the equations you get \[x - 7 = 15\]\[x = 22\] So the solution is \[(22,7)\]

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

could you help me with this one too? 4x-7y=3 x-7y= -15

OpenStudy (anonymous):

Multiply one of the equations by -1 so you can eliminate the 7y's

OpenStudy (anonymous):

It doesnt matter which one

OpenStudy (anonymous):

so then: (4x-7y=3) -1= -4x+7y= -3 x-7y= -15 and that would be... 3x= -18?

OpenStudy (anonymous):

-3x = -18.. you where close

OpenStudy (anonymous):

Then when you plug x back into one of the equations be sure to use the equations that where given

OpenStudy (anonymous):

Not the equation you multiplied -1 by

OpenStudy (anonymous):

really? yesss. so even if the x's are negative i still add them? never actually subtract? because what i did was: 4x-x= 3x and what i was supposed to do was: -4x+x= =3

OpenStudy (anonymous):

I should say. not the equation you got when you multiplied it by -1

OpenStudy (anonymous):

well if you subtract them you would get -4x - x = -5x

OpenStudy (anonymous):

ya i know, thats what i keep forgetting about. like i actually subtracted them instead of adding them. is there a time when i actually COULD subtract them? so then to solve for y it would be: 6-7y= -15 -6 -6 -7y/-7y= -21/-7y y= 3?

OpenStudy (anonymous):

so you would get x to be 6.. then plug x back into one of the original equations and you get \[6 - 7y = -15\]\[-7y = -21\]\[y = 3\] So the solution is (6,3).. and to check this plug in both numbers into one of the equations \[6 - 7(3) = -15\]\[6 - 21 = -15 \space {\checkmark}\] So its correct :)

OpenStudy (anonymous):

YAY!! thank you sooo much :D i really appreciate it :)

OpenStudy (anonymous):

No problemm